# A comparison of simulation tools for Muon-Induced X-ray Emission (MIXE) in thin films: a study case with lithium batteries

### v1

@@ Abstract
We present a comparative study of three Monte Carlo simulation frameworks—SRIM, GEANT4, and PHITS—for modeling the transport, stopping, and atomic cascade of negative muons in micrometer-scale, multilayer systems relevant to Muon-Induced X-ray Emission (MIXE) experiments at the Paul Scherrer Institute (PSI). Using a lithium-ion battery as a benchmark target, simulated implantation profiles are compared with experimental data from the GIANT spectrometer. All three codes reproduce the overall muon depth distributions with good consistency, even across sharp density contrasts. SRIM provides reliable implantation estimates for compact geometries, whereas PHITS reproduces GEANT4 results with comparable accuracy and additionally generates muonic X-ray spectra. These spectra, however, exhibit a systematic energy offset in the K-line transitions of medium- and high-Z elements relative to theoretical and experimental values. Despite this bias, PHITS accurately captures relative intensities and spectral shapes, enabling element-specific line identification. The results demonstrate that SRIM and PHITS constitute practical tools for rapid estimation of muon implantation and stopping profiles, and that PHITS holds strong potential for predictive MIXE spectroscopy once its transition-energy bias is corrected.
@@ 1 Introduction, MIXE at Paul Scherrer Institute
Muon-Induced X-ray Emission (MIXE) has long been employed to probe atomic-scale properties such as nuclear size, shape and charge distribution [anderson1969, muX2021]. The field originated in the 1960s–1970s through precision muonic-atom spectroscopy, where measurements of transition energies revealed fine nuclear-structure effects beyond the reach of conventional X-ray methods. In the 1980s, pioneering work at the Swiss Institute for Nuclear Research (SIN)—now the Paul Scherrer Institute (PSI)—extended MIXE to non-destructive elemental analysis of cultural and geological samples [MIXE_PSI_1981, MIXE_PSI_1981_2]. With PSI’s high-intensity continuous muon beams, generated at the High-Intensity-Proton-Accelerator-Facility HIPA [HIPA_1], the Swiss Muon Source (SµS) now enables routine MIXE studies of bulk specimens using the GIANT HPGe array  [GIANT_2023] (Fig. 1). Recent efforts within the Swiss National Science Foundation (SNSF) interdisciplinary research program DEEPµ(SNSF Sinergia grant 193691) have demonstrated depth-resolved, isotope-specific analysis of materials ranging from archaeological artifacts to energy-storage devices, motivating the development of reliable simulation tools for experiment design and data interpretation.
In the MIXE technique at PSI, negative muons from the E1 beamline are implanted in a sample with tuneable momenta —from 15 to 60 MeV/c —corresponding to a penetration range up to $\mathrm{3~g/cm^{2}}$. Momenta of up to 125 MeV/c (58 MeV kinetic energy), and therefore a deeper penetration of the sample, can be achieved in the future by modifying the E1 beamline. This specific beamline delivers muons originating from in-flight pion decay extracted from the thick (E) graphite target [HIPA_1], in a long superconducting magnet. Momenta between 60 to 125 MeV/c can be selected, allowing to probe from millimeters to about 10 centimeters in aluminium.
@@
At the end of the stopping process in the sample, negative muons are captured in the Coulomb field of the positive charge of the target nucleus, creating a so-called muonic atom [anderson1969, MIXE_PSI_1981, muX2021]. This capture will leave the muon-atom system in an excited state with the principal quantum number $n\approx 14$ [KNECHT_2020], which subsequently decays to the ground state within a picosecond resulting in the emission of characteristic X-rays . This emission process can be compared to electron-induced X-ray fluorescence, but at 207 times higher energies due to the larger muon mass. Fig. 2 depicts the 3 main processes of MIXE: muon implantation, muon capture, and muonic cascade.
@@
An approximation of the X-ray energies released in this process may be computed using the Dirac equation, like for electronic transitions, but with the mass of the muon, 207 times larger than the electron. The energy range of muon-induced X-ray fluorescence is therefore around two orders of magnitude larger than that of electron-induced fluorescence. Elements heavier than vanadium all display K1-L3 emission lines with energies greater than 1 MeV. Such energetic photons, unlike soft X-rays induced by electronic cascade, can traverse long distances in dense material, such as a metallic battery pouch [Izumi_2018, Izumi_2020]. Computer codes such as Akylas[Akylas_code] or MUDIRAC[MUDIRAC_2021] can produce an extended database of typical X-ray photons, taking into account both isotopic and orbital differences. PHITS [mandatoryPHITS, phits_333, phits_benchmark] and GEANT4 [mandatoryG4, cataldo2023G4MIXE] can compute these typical transitions and output a tentative muonic X-ray photon spectrum.
Once the captured negative muons reach the ground state, they either decay or undergo nuclear capture, especially in high-Z elements. This capture may additionally lead to gamma and/or neutron emission, either promptly or when the transmuted isotopes decay. These decays can be recorded for subsequent analysis and cross-verification of the isotope identified by MIXE.
Heavy elements (Z>>25) display line splittings due to an isotopic shift on the order of several keV, allowing for easy non-destructive identification of isotopes. Using the ratio of $K$ X-ray line emissions of materials bombarded by negative muons, it is possible to assess the isotopic composition of the sample. This technique was successfully employed for the non-destructive determination of the geographic origin of lead (Pb) artifacts, as its isotopic composition varies with mining locations[ninomiya2019_lead_isotopes]. In this case, the $K$ energy lines are found to be around 6 MeV, minimizing self-shielding, thereby facilitating accurate analysis at greater depths.
@@
Figure 2: Steps to be modeled in PHITS: (1) negative muon transport from the beamline to the sample; (2) capture within the Coulomb field of a target atom; (3) muonic cascade with simultaneous emission of muonic X-rays. Image extracted from the work of Quérel et al. [edouard]
@@
The sensitivity of MIXE to light isotopes (Na, Li) was first demonstrated in Japan [Izumi_2018, Izumi_2020], and this result enabled a team from the Swiss Federal Laboratories for Material Sciences and Technology (Empa) to begin applying MIXE for in-situ, live, following of electrochemical processes inside battery packs [edouard]. These experimental setups and data will be implemented and compared to numerical models described below.
@@ 2.1 Experimental geometry: the GIANT setup
The GIANT setup (Fig. 1) consists of an aluminium frame hosting up to 30 high-purity germanium (HPGe) detectors, continuously cooled by a liquid nitrogen circulation system. Samples are analyzed in air, positioned 10–20 cm from the beamport window depending on detector heads placements. Negative muons from the High-Intensity Proton Accelerator (HIPA) [HIPA_1] are transported by the E1 beamline and focused by a quadrupole triplet onto a plastic scintillator tagger (BC-400, 21×22 mm2, 200 µm thick). The light output is read by Advansid NUV-series SiPMs, providing $\sim$0.4 ns timing resolution. After tagging, muons cross a 10 µm titanium window into air before reaching the sample. Multiple scattering in these entrance materials slightly broadens the beam and increases angular divergence at the target.
@@ 2.2 Simulation tools and configuration
Since SRIM does not natively support muons, pseudo-muons were modeled as protons with rescaled mass $m=105.7/931.5\simeq 0.113$ amu. Because SRIM’s stopping powers scale with $Z^{2}$ and rely on Coulomb scattering, this approximation reproduces muon energy loss reasonably well while avoiding charge-exchange artifacts. Despite its limitations, SRIM (and TRIM.SP) has been successfully applied to positive-muon transport in thin films [TRIM_Elvezio], and here serves as a lightweight benchmark for MIXE energies. Importantly, SRIM offers no possibility to simulate realistic beam optics: only monoenergetic, pencil-like beams incident on semi-infinite targets can be used.
@@
Recent improvements in the PHITS muon-interaction module [PHITS_muon] motivated its benchmarking against GEANT4, the reference toolkit for particle transport. Muon capture cross sections were taken from Suzuki et al. [CaptureXS_Suzuki_1987], and the ensuing muonic X-ray cascade was simulated using the Akylas–Vogel routine [Akylas_code]. Each code modeled the same simplified geometry (Fig. 3), including the tagger, vacuum window, 10 cm air gap, and eight battery layers with densities and thicknesses listed in Table 1. Beam momentum spreads of $\pm$0.5% around nominal momenta were implemented to reflect the E1 beam characteristics. A nominally pencil-like muon beam was generated in both codes; however, the multiple scattering induced by the muon tagger results in a significantly broadened beam spot after 10 cm of air gap, effectively defining the incident profile on the sample.
@@
Figure 3: Left: PHITS geometry showing the muon beam (from left) through the tagger (yellow) and titanium window (green) before reaching the sample (red) 14 cm downstream. Right: Zoom on the multilayer battery stack. Material parameters are listed in Table 1.
Table 1: Thickness and density of the battery layers. The pouch is a polyamide–Al–polyethylene composite.
@@ 2.3 Simulation scope and precision
PHITS offers multiple physics options for transport accuracy; the key cards used are listed in Appendix A. A sensitivity study showed that the nspred (Coulomb diffusion) flag alters implantation depth by less than 1%, indicating minimal angular-straggling influence under the present conditions. The GEANT4 application uses physics library QGSP_BIC_HP_EMZ and G4 material database.
@@
3 Results and discussion
3.1 Negative muon implantation depths
A total of 56 muon energy groups were analyzed for their penetration and implantation depths within the modeled battery cell. The momentum limits were set by the lowest energy at which muons reach the target—many being stopped in the titanium window—and the highest energy where muons traverse the full stack without stopping. For GEANT4 and PHITS, momenta between 20 MeV/c and 30 MeV/c were simulated with up to $2\times 10^{6}$ primaries, after preliminary $10^{4}$-event runs to set boundaries of our energy scan range. SRIM results were obtained from $10^{4}$ pseudo-muons. The mean implantation depths computed by each code are summarized in Tab. LABEL:tab:depths.
Table 2: Depths in µm of mean muon implantation in the battery, as computed by GEANT4 (reference), PHITS, and SRIM. Muons stopping in air before the battery are labeled "air".
@@ 3.1 Negative muon implantation depths
While GEANT4 and PHITS results are nearly identical, SRIM exhibits significant deviations, especially when large air volumes precede the thin-film stack. To isolate the origin of this discrepancy, the range of pseudo-muons in each battery layer was recalculated with SRIM and PHITS (Tab. 3).
Table 3: Ranges in µm of muons in individual battery layers at various momenta $P$, computed by PHITS (positive muons) and SRIM (pseudo-muons).
@@
For isolated layers, SRIM and PHITS agree within 5%. Additional tests on upstream materials (scintillator, titanium window, and various gases) confirmed this consistency (Tab. 4).
Table 4: Ranges in µm of muons in components along the beam path, computed by PHITS (positive muons) and SRIM (pseudo-muons).
Layer He (m) Air (m) CO2 (m) BC400 (mm) Ti (µm)
P (MeV/c) 20 30 20 30 20 30 20 30 20 30
PHITS 1.87 8.23 0.32 1.37 0.21 0.88 0.33 1.40 117 478
SRIM 1.92 8.25 0.326 1.38 0.214 0.90 0.329 1.40 114 467
Diff (%) 2.67 0.24 1.88 0.73 1.90 2.27 0.30 0.00 2.56 2.30

### v2

@@ Abstract
We present a comparative study of three Monte Carlo simulation frameworks—SRIM, GEANT4, and PHITS—for modeling the transport, stopping, and atomic cascade of negative muons in micrometer-scale, multilayer systems relevant to Muon-Induced X-ray Emission (MIXE) experiments at the Paul Scherrer Institute (PSI). Using a lithium-ion battery as a benchmark target, simulated stopping depth profiles are compared with experimental data from the GIANT spectrometer. All three codes reproduce the overall muon depth distributions with good consistency, even across sharp density contrasts. SRIM provides reliable stopping depth estimates for compact geometries, whereas PHITS reproduces GEANT4 results with comparable accuracy and additionally generates muonic X-ray spectra. These spectra, however, exhibit a systematic energy offset in the K-line transitions of medium- and high-Z elements relative to theoretical and experimental values. Despite this bias, PHITS accurately captures relative intensities and spectral shapes, enabling element-specific line identification. The results demonstrate that SRIM and PHITS constitute practical tools for rapid estimation of muon stopping depth and stopping profiles, and that PHITS holds strong potential for predictive MIXE spectroscopy once its transition-energy bias is corrected.
@@ 1 Introduction, MIXE at Paul Scherrer Institute
Muon-Induced X-ray Emission (MIXE) has long been employed to probe atomic-scale properties such as nuclear size, shape and charge distribution [anderson1969, muX2021]. The field originated in the 1960s–1970s through precision muonic-atom spectroscopy, where measurements of transition energies revealed fine nuclear-structure effects beyond the reach of conventional X-ray methods. In the 1980s, pioneering work at the Swiss Institute for Nuclear Research (SIN)—now the Paul Scherrer Institute (PSI)—extended MIXE to non-destructive elemental analysis of cultural and geological samples [MIXE_PSI_1981, MIXE_PSI_1981_2]. With PSI’s high-intensity continuous muon beams, generated at the High-Intensity-Proton-Accelerator-Facility HIPA [HIPA_1], the Swiss Muon Source (SµS) now enables routine MIXE studies of bulk specimens using the GIANT HPGe array [GIANT_2023] (Fig. 1). Recent work conducted within the framework of the Swiss National Science Foundation (SNSF) interdisciplinary research program DEEPµ (SNSF Sinergia grant 193691) has enabled the application of MIXE across a broad range of fields—similarly to initiatives undertaken at other facilities—including depth-resolved, isotope-specific investigations of archaeological artifacts[cataldo2022novel, biswas2023, hillier2025coin], scientifically significant objects[rossini2023meteorite, terada2014meteorite, ninomiya2019_lead_isotopes], and energy-storage devices[Izumi_2020, edouard]. These developments have further emphasized the need for robust simulation tools to support experimental design and data interpretation [cataldo2025MonteCarlo].
In the following, denotes a negative muon (-) unless stated otherwise. MIXE relies on - atomic capture and the ensuing muonic-atom cascade; positive muons (+) do not form bound muonic atoms and are therefore used primarily for SR-type measurements rather than muonic X-ray emission.
@@
For layered or encapsulated systems, MIXE is complementary to conventional ion-beam techniques such as Rutherford Backscattering Spectrometry (RBS) and Nuclear Reaction Analysis (NRA), as well as synchrotron-based X-ray methods. The high energy of muonic X-rays (typically 0.1 up to 10 MeV) reduces self-absorption compared to conventional XRF and enables analysis through thick overlayers or shells. Unlike Neutron Activation Analysis (NAA), the sample is not activated and can be returned shortly after measurement. A limitation is the cm-scale beam spot, implying limited lateral resolution compared with micro-focused ion or synchrotron probes; MIXE is therefore best viewed as a complementary tool for non-destructive, depth-resolved bulk analysis of complex objects.
@@
At the end of the stopping process in the sample, muons are captured in the Coulomb field of the positive charge of the target nucleus, creating a so-called muonic atom [anderson1969, muX2021, MIXE_PSI_1981]. This capture will leave the muon-atom system in an excited state with the principal quantum number $n\approx 14$ [KNECHT_2020], which subsequently decays to the ground state within a picosecond resulting in the emission of characteristic X-rays. This emission process can be compared to electronic X-ray fluorescence, but at two orders of magnitude higher energies due to the 207 times larger muon mass. The physics of muonic atoms and cascade processes is extensively described elsewhere [hillier2022muon, amato2024Book]. Fig. 2 depicts the 3 main processes of MIXE: muon stopping, muon capture, and muonic cascade.
@@
Computer codes such as Akylas [Akylas_code] or MuDirac [MuDirac_2021] can produce an extended database of typical X-ray photons originating muonic cascade, taking into account both isotopic and orbital differences. PHITS [mandatoryPHITS, phits_333, phits_benchmark] and GEANT4 [mandatoryG4, cataldo2023G4MIXE] can compute these typical transitions and output a tentative muonic X-ray photon spectrum.
@@
Figure 2: Steps to be modeled in PHITS: (1) muon transport from the beamline to the sample; (2) capture within the Coulomb field of a target atom; (3) muonic cascade with simultaneous emission of muonic X-rays. Image extracted from the work of Quérel et al. [edouard]
In the MIXE technique at PSI, muons delivered by the E1 beamline are implanted into the sample with adjustable momenta between 15 and 60 MeV/c, corresponding to a maximum penetration of about $\mathrm{3~g/cm^{2}}$. An upgrade of the E1 beamline is planned to extend this range up to 125 MeV/c, enabling substantially deeper stopping depths. This beamline transports muons from in-flight pion decay produced in the thick graphite Target E station at HIPA [HIPA_1] through a long superconducting magnet. In the upgraded configuration, momenta between 60 and 125 MeV/c will be selectable, corresponding, for example, to penetration depths from the millimeter scale up to approximately 10 cm in aluminium.
@@
The sensitivity of MIXE to light isotopes (Na, Li) was first demonstrated in Japan [Izumi_2020, Izumi_2018], and this result enabled a team from the Swiss Federal Laboratories for Material Sciences and Technology (Empa) to begin applying MIXE for in situ following of electrochemical processes inside battery cells [edouard]. This experimental setup and data and will be implemented in the simulation frameworks and compared to numerical models described below.
The objective of this work is therefore to identify simulation tools that are both reliable and practical for predicting muon stopping profiles and muonic X-ray spectra in layered materials, with particular emphasis on usability for experiment preparation. The benchmarking strategy adopted here (transport/stopping validation and spectral-model validation) is not specific to lithium-ion batteries and can be transferred to other heterogeneous layered or encapsulated systems such as cultural-heritage coatings/corrosion layers, geological/extraterrestrial specimens, and solid-state battery architectures. Using a multilayer lithium-ion battery as a representative heterogeneous target with demanding simulation requirements, we benchmark SRIM, GEANT4, and PHITS against each other and against experimental measurements. The study aims to determine the level of accuracy that can be expected from each code and to provide guidance for future users selecting simulation tools for MIXE-based investigations. We emphasize that the present benchmark focuses on transport/stopping accuracy and on the internal consistency of cascade-generated spectra; it does not attempt a full detector-response model or an absolute-efficiency validated prediction of experimental count rates.
@@ 2.1 Experimental geometry: the GIANT setup
The GIANT setup (Fig. 1) consists of an aluminium frame hosting up to 30 high-purity germanium (HPGe) detectors, continuously cooled by a liquid nitrogen circulation system. Samples are analyzed in air, positioned 10–20 cm from the beamport window depending on detector placements. Muons from the High-Intensity Proton Accelerator (HIPA) [HIPA_1] are transported by the E1 beamline and focused by a quadrupole triplet onto a plastic scintillator tagger (BC-400, 21×22 mm2, 200 µm thick). The light output is read by Advansid NUV-series SiPMs, providing $\sim$0.4 ns timing resolution. After tagging, muons cross a 10 µm titanium window into air before reaching the sample. Multiple scattering in these entrance materials slightly broadens the beam and increases angular divergence at the target.
@@ 2.2 Simulation tools and configuration
Since SRIM does not natively support muons, pseudo-muons were modeled as protons with rescaled mass $m=105.7/931.5\simeq 0.113$ amu. Because SRIM’s stopping powers scale with $Z^{2}$ and rely on Coulomb scattering, this approximation reproduces muon energy loss reasonably well while avoiding charge-exchange artifacts. Despite its limitations, SRIM (and TRIM.SP [Biersack1984], an ion transport and sputtering code) has been successfully applied to positive-muon transport in thin films [TRIM_Elvezio], and here serves as a lightweight benchmark for MIXE energies. Importantly, SRIM offers no possibility to simulate realistic beam optics: only monoenergetic, pencil-like beams incident on semi-infinite targets can be used.
@@
Recent improvements in the PHITS muon-interaction module [PHITS_muon] motivated its benchmarking against GEANT4, the reference toolkit for particle transport. Muon capture cross sections were taken from Suzuki et al. [CaptureXS_Suzuki_1987], and the ensuing muonic X-ray cascade was simulated using the Akylas–Vogel routine [Akylas_code]. Each code modeled the same simplified geometry (Fig. 3), including the tagger, vacuum window, 10 cm air gap, and eight battery layers with densities and thicknesses listed in Tab. 1. Beam momentum spreads of $\pm$0.5% around nominal momenta were implemented to reflect the E1 beam characteristics. A nominally pencil-like muon beam was generated in both codes; however, the multiple scattering induced by the muon tagger results in a significantly broadened beam spot after 10 cm of air gap, effectively defining the incident profile on the sample.
@@
Figure 3: Left: PHITS geometry showing the muon beam (from left) through the tagger (yellow) and titanium window (green) before reaching the sample (red) 14 cm downstream. Right: Zoom on the multilayer battery stack. Material parameters are listed in Tab. 1.
Table 1: Thickness and density of the battery layers. The pouch is a polyamide–Al–polyethylene composite. Here, “Whatman” denotes a glass-fiber separator, whereas “Celgard” denotes a thin microporous polymer separator film.
@@ 2.3 Simulation scope and precision
PHITS offers multiple physics options for transport accuracy; the key cards used are listed in Appendix A. A sensitivity study showed that the nspred (Coulomb diffusion) flag alters stopping depth by less than 1%, indicating minimal angular-straggling influence under the present conditions. The GEANT4 application uses physics library QGSP_BIC_HP_EMZ and G4 material database.
Statistical uncertainties on simulated observables were evaluated from counting statistics within each histogram bin assuming Poisson behavior and propagated to derived quantities (mean stopping depth, peak intensities, and fitted Gaussian areas). In the present simulations ($\sim 10^{6}$ primaries), the standard error of the mean (SEM) on the mean stopping depth (MSD) is sub-µm and does not limit the GEANT4/PHITS comparison. Additional systematic contributions may arise from histogram binning choices, detector response assumptions used for peak integration, from imperfect knowledge of the experimental geometry (layer thicknesses, material densities, air gaps and alignment), and detector efficiency. Manufacturer tolerances for the benchmark cell were not available and these geometry-related uncertainties are therefore not quantified here; they should be regarded as a potential dominant contribution when comparing to experimental data. Output stopping depth profiles were averaged over the final 5% of trajectories to derive the mean range and standard deviation. The results are compared quantitatively in the next section.
3 Results and discussion
3.1 Muon stopping depths
3.1.1 Momentum scan and stopping-depth comparison
A total of 56 muon momenta were investigated to determine the corresponding penetration and stopping depths within the modeled battery cell. The lower momentum bound was defined by the minimum value at which muons reach the target, as a significant fraction is otherwise stopped in the titanium window, while the upper bound corresponds to momenta for which muons traverse the entire stack without stopping.
For both GEANT4 and PHITS, momenta between 20 MeV/c and 30 MeV/c were simulated with up to $2\times 10^{6}$ primary particles, following preliminary runs of $10^{4}$ events to establish the limits of the momentum scan. SRIM results were obtained using $10^{4}$ pseudo-muons. The MSD calculated with each code is summarized in Tab. 2.
3.1.2 Uncertainty estimates
To quantify the statistical reliability of the simulated stopping depths, the MSD was computed from the stopping-position distribution and its spread (range straggling) defined as:
@@ 3.1.2 Uncertainty estimates
where $x_{i}$ denotes the depth coordinate and $w_{i}$ the number of stopped muons in bin index $i$. The statistical uncertainty on the MSD, the SEM, was estimated as
[…]
with $N_{\mathrm{eff}}=\sum_{i}w_{i}$ the effective number of stopped muons. As mentioned, the resulting statistical uncertainties are below 0.1 µm and therefore negligible compared to the intrinsic stopping depth spread , which is typically several to a few tens of micrometers, depending on material and momentum, as compiled in Tab.4. When comparing to experimental spectra, additional systematic contributions (geometry, detector efficiency, background subtraction, and low-energy attenuation) can dominate and are discussed separately in Sec. 3.2.
Table 2: Mean stopping depths (in µm) in the battery, computed by GEANT4 (baseline), PHITS, and SRIM. The SEM on the Monte Carlo mean depth is <0.1<0.1 µm for GEANT4 and PHITS at all momenta for the present statistics.
@@
While GEANT4 and PHITS results are nearly identical, SRIM exhibits significant discrepancies, especially when large air volumes precede the thin-film stack. To isolate the origin of this discrepancy, the range of pseudo-muons in each battery layer was recalculated with SRIM and PHITS (Tab. 3).
3.1.3 SRIM limitations in low-density media
Previous studies[cataldo2023Air] have reported that SRIM range predictions can be improved when extended low-density regions are replaced by compressed-density equivalents, effectively reducing numerical precision limitations associated with large path lengths. While this approach can reproduce realistic stopping positions, it alters the physical geometry and therefore was not adopted here, since the goal of this work is a direct comparison of transport codes under identical and physically faithful conditions. We note that such density scaling may be useful for rapid range estimation but does not remove intrinsic precision limitations when long trajectories must be explicitly tracked.
Table 3: Mean stopping depth in µm of muons in individual battery layers at various momenta $P$, computed by PHITS (baseline) and SRIM. The SEM on the Monte Carlo mean stopping depth is <0.1<0.1 µm for PHITS at all momenta for the present statistics.
@@ 3.1.3 SRIM limitations in low-density media
For isolated layers, SRIM and PHITS agree within 5%. Additional tests on upstream materials (scintillator, titanium window, and various gases) confirmed this consistency (Tab. 4). However, we note a larger SRIM-computed spread in titanium and scintillator. An important difference between PHITS and SRIM algorithms arises from the treatment of negative muon nuclear capture. SRIM/TRIM does not include muon capture physics and therefore describes purely electromagnetic slowing down and stopping.
3.1.4 Effect of muon capture in upstream materials
In PHITS, when negative muon capture is enabled, a fraction of the muons are absorbed by the nucleus shortly after atomic capture, modifying the spatial distribution of disappearance points. As muonic capture probability increases with $Z$, we additionally repeated the PHITS runs with muon capture imucap disabled in the last column of Tab. 4 to isolate purely electromagnetic stopping and assess its impact on the disappearance-depth distribution.
Table 4: Mean stopping depth and stopping-depth spread (, i.e. range straggling) for selected upstream materials as computed by SRIM and PHITS, with and without the negative-muon capture flag imucap. The SEM on the Monte Carlo mean depth is <0.1<0.1 µm for PHITS at all momenta for the present statistics.
Layer P (MeV/c) PHITS (w. mu) SRIM PHITS (w/o mu)
He (cm) 20 189 7.33 192 7.37 189 7.27
He (cm) 30 829 2.80 825 3.20 830 2.87
Air (cm) 20 32.1 1.18 34.0 1.91 32.3 1.55
Air (cm) 30 137 5.37 138 6.43 135 5.05
CO2 (cm) 20 21.0 0.73 21.4 1.01 20.6 0.89
CO2 (cm) 30 88.0 6.46 90.1 4.17 87.2 5.75
BC400 (µm) 20 325 1.74 330 16.0 325 1.78
BC400 (mm) 30 1.40 0.06 1.42 0.06 1.40 0.06
Ti (µm) 20 117 2.80 114 8.68 117 2.82
Ti (µm) 30 478 9.43 465 33.0 476 9.43